# Computing the Intrinsic Delaunay Triangulation of a Closed Polyhedral Surface

### v1

@@ Abstract
keywords:
Polyhedral surface, intrinsic Delaunay triangulation, algorithmic complexity
@@ 1 Introduction
In one of its simplest forms a triangulation is a finite collection of disjoint triangles in the Euclidean plane, together with a partial matching of the sides of the triangles such that any two matched sides have the same length (Figure 1). This simple representation appears under different names in the literature (intrinsic triangulation [sharp2021geometry, sharp2019navigating], portalgon [portalgons]). Cutting out the triangles from the plane and identifying the matched sides isometrically, respecting the orientations of the triangles, provides a (compact, orientable) polyhedral surface. This surface is closed if, in addition, it is connected and without boundary.
In this paper we consider the Delaunay triangulation of a closed polyhedral surface whose vertices are the singularities (the points surrounded by an angle distinct from $2\pi$) of the surface (except for flat tori, see below). It is generically unique. Our main contribution is an algorithm to compute it from any triangulation of the surface, whose time complexity is polynomial in the number of triangles and in the logarithm of the aspect ratio of the input triangulation. Our second contribution is a lower bound showing that the dependency in the logarithm of the aspect ratio is unavoidable in our model of computation.
@@ 1.1 Related works
Polyhedral surfaces can also be obtained from meshes, flat triangles in $\mathbb{R}^{3}$ glued along their edges. Moreover, every mesh defines a triangulation of its surface. Yet triangulations are more general than meshes: most triangulations cannot be obtained from a mesh. Some recent algorithms advantageously operate on triangulations of polyhedral surfaces without reference to a mesh [sharp2020you, takayama2022compatible, liu2023surface]. In this context the adjective “intrinsic” is sometimes placed before the name “triangulation” to make the distinction with the particular triangulations arising from a mesh. In the mathematical community, a prominent example is that of a translation surface [masur2006ergodic, zorich2006flat, hubert2006introduction, gutkin2000affine], which arises naturally in the study of billiards in rational polygons.
@@
Triangulations are so general that not all of them are suitable for computation, compared to meshes. Prominently, a fundamental problem on polyhedral surfaces is to compute the distance or a report a shortest path between two points. On a mesh, shortest paths can be computed in time polynomial in the number of triangles. For example, an algorithm of Mitchell, Mount, and Papadimitriou [mitchell1987discrete] propagates waves along the surface, starting from the source, in a discrete manner. See also Chen and Han [chen1996shortest]. On a generic triangulation however (not arising from a mesh), the number of times a shortest path visits a triangle is not bounded by any function of the number of triangles, as noted for example almost 20 years ago by Erickson [bworld]. Recently, Löffler, Ophelders, Staals, and Silveira [portalgons] coined the term happiness of a triangulation, for the maximum number of times a shortest path visits a triangle. They adapted the single-source shortest paths algorithm of Mitchell, Mount, and Papadimitriou [mitchell1987discrete] from meshes to triangulations, whose time complexity now depends on the happiness of the triangulation (it is more efficient on triangulations of low happiness).
@@
This raises the problem of replacing any given triangulation by another triangulation of the same surface whose happiness is “low”. Among the many remeshing algorithms [heckbert1997survey, shewchuk1997delaunay, ruppert1995delaunay, shewchuk2002delaunay], only few have been ported to the general context of intrinsic triangulations [sharp2021geometry], and the only solution we are aware of that compares to our main result (Theorem 1.1 below), by Löffler, Ophelders, Staals, and Silveira [portalgons, Section 5], is restricted to particular inputs whose surfaces are all homeomorphic to an annulus; we will use it as a black box. Importantly, the same authors also showed that Delaunay triangulations have bounded happiness [portalgons, Section 4.2].
@@
Delaunay triangulations are classical objects of computational geometry [de2000computational, fortune2017voronoi, aurenhammer2013voronoi], closely related to shortest paths. While mostly known in the plane, they generalize to closed polyhedral surfaces, see for example the depiction of Bobenko and Springborn [bobenko2007discrete]. To compute a Delaunay triangulation from an arbitrary intrinsic triangulation there are, to our knowledge, only two approaches, and neither compares to our main result (Theorem 1.1). One approach computes a Voronoi diagram with a suitably adapted multiple-source shortest path algorithm, and then derives from it a Delaunay tessellation, see for example Mount [mount1985voronoi], Liu, Chen, and Tang [liu2010construction], and Liu, Xu, Fan, and He [liu2015efficient]. Another approach starts from an initial triangulation and flips its edges until it reaches a Delaunay triangulation; it was proved to terminate by Indermitte, Liebling, Troyanov, and Clémençon [bobenko2007discrete, indermitte2001voronoi].
@@ 1.2 Our results
In order to state our results precisely, it now matters to make the distinction between a triangulation and the data structure representing it, and to allow for more general polygons than triangles. Following Löffler, Ophelders, Staals, and Silveira [portalgons], we call portalgon the collection $T$ of polygons in the Euclidean plane and the partial matching of their sides. We denote by $\mathcal{S}(T)$ the associated polyhedral surface. We say that the portalgon $T$ is triangular if all polygons are triangles. The sides of the polygons, once identified, correspond to graph $T^{1}$ embedded on $\mathcal{S}(T)$: it is this graph $T^{1}$ that we call triangulation if $T$ is triangular, and we call $T^{1}$ a tessellation in general.
We consider, on a closed polyhedral surface $S$, the unique Delaunay tessellation $\mathcal{D}$ of $S$ whose vertices are exactly the singularities of $S$, with a single very special exception: if $S$ has no singularity, then $S$ is a flat torus and we let $\mathcal{D}$ be any of the Delaunay tessellations of $S$ that have exactly one vertex, for one can be mapped to the other via an orientation-preserving isometry of $S$ anyway. In any case, we say that $\mathcal{D}$ is the Delaunay tessellation of $S$, in a slight abuse. [^2] Given a set $V$ of points of the surface, finite, non-empty, and containing all the singularities, our results easily extend to Delaunay triangulations whose vertex set is $V$, but this is incidental to us. It is “generically” a triangulation, but not always. If not, then triangulating the faces of $\mathcal{D}$ along vertex-to-vertex arcs provides a Delaunay triangulation. The aspect ratio of a triangular portalgon $T$ is the maximum side length of a triangle of $T$ divided by the smallest height of a triangle of $T$ (possibly another triangle). Our main contribution is:
@@
We analyze our algorithms within the real RAM model of computation described by Erickson, van der Hoog, and Miltzow [erickson2022smoothing]. It is an extension of the standard integer word RAM, with an additional memory array storing reals, and with additional instructions. On such a machine, we represent each polygon of a portalgon $T$ by the list of its vertices, and each vertex is by its two coordinates, stored in the memory array dedicated to reals. So displacing the polygons in the plane provides different representations of $T$. When modifying a portalgon $T$, we actually modify our representation of $T$, using elementary operations that are easily seen to be achievable by a real RAM.
@@
Proof 1.6 (Proof of Theorem 1.1).
Proposition 1.4 computes in $O(n\log^{2}(n)\cdot\log^{2}(r))$ time a portalgon $T^{\prime}$ of $O(n\cdot\log(r))$ triangles, whose happiness is $O(n\log(n)\cdot\log^{2}(r))$. Proposition 1.5 then computes the portalgon of the Delaunay tessellation from $T^{\prime}$ in $O(n^{3}\log^{2}(n)\cdot\log^{4}(r))$ time.
@@ 2 Preliminaries
We use without review standard notions of graph theory and low dimensional topology and geometry, referring to textbooks for details [diestel, armstrong2013basic, s-ctcgt-93, do1992riemannian]. We only mention that on a surface $S$, a path $p:[0,1]\to S$ is simple if its restriction to the interval $(0,1)$ is injective, in which case the image of $(0,1)$ by $p$ is the relative interior of $p$. We denote by $\ell(p)$ the length of a geodesic path $p$. Throughout the paper, logarithms are in base two.
The definition of Delaunay tessellation given by Bobenko and Springborn [bobenko2007discrete, Section 2] is not used in the core of the paper, but only in in Appendix J for proving Proposition 1.5. We collect details on this definition in Appendix I for completeness.
@@ 2.1 Portalgons, tessellations, and polyhedral surfaces
A portalgon $T$ is a disjoint collection of oriented polygons in the Euclidean plane, together with a partial matching of the sides of the polygons such that any two matched sides have the same length. It is triangular if all polygons are triangles. See Figure 1. Any subset of the polygons defines a sub-portalgon $T^{\prime}$ of $T$: two sides of polygons are matched in $T^{\prime}$ if and only if they are matched in $T$. In a portalgon $T$, identifying the matched sides, isometrically, and respecting the orientations of the polygons, provides the surface of $T$, denoted $\mathcal{S}(T)$; it is a 2-dimensional Riemannian manifold whose metric may have singularities. The sides of the polygons of $T$ correspond to a graph $T^{1}$ embedded on $\mathcal{S}(T)$, the 1-skeleton of $T$.
@@
Figure 1: (Left) A triangular portalgon $T$: two triangles in the Euclidean plane, with two sides matched in red. (Right) The surface $\mathcal{S}(T)$, and the 1-skeleton $T^{1}$.
@@ 2.2 Aspect ratio, systole, happiness, and segment-happiness
The aspect ratio of a triangular portalgon $T$ is the maximum side length of a triangle of $T$ divided by the smallest height of a triangle of $T$ (possibly another triangle). Note that the aspect ratio is always greater than or equal to $\sqrt{3}/2>1$, because the maximum side length of a triangle is always greater than or equal to $\sqrt{3}/2$ times its smallest height.
@@
The happiness of a portalgon $T$ is the maximum number of times a shortest path in $\mathcal{S}(T)$ visits the image of a polygon of $T$, maximized over all the shortest paths of $\mathcal{S}(T)$ and all the polygons of $T$ (see [portalgons, Section 3]). We introduce a variation, more suitable to our needs. In a polyhedral surface $S$, a segment is a simple geodesic path $e$ whose relative interior is disjoint from any singularity of $S$. The segment-happiness of $e$ in $S$, denoted $h_{S}(e)$, is the maximum number of intersections between $e$ and a shortest path of $S$, maximized over all the shortest paths of $S$. The segment-happiness of a portalgon $T$ is then the maximum segment-happiness $h_{\mathcal{S}(T)}(e)$, maximized over the edges $e$ of its 1-skeleton $T^{1}$. A priori, the segment-happiness of a portalgon $T$ differs from the happiness of $T$. Indeed a path in $\mathcal{S}(T)$ may visit many times a face of $T^{1}$ without intersecting any edge of $T^{1}$ more than once, if the face has high degree. However, if $T$ is triangular, then the happiness and the segment-happiness of $T$ do not differ by more than a constant factor.
@@ 3 Tubes and bifaces
In this section we focus on particular triangular portalgons. See Figure 2. A tube is a triangular portalgon $X$ whose surface $\mathcal{S}(X)$ is homeomorphic to an annulus and has no singularity in its interior, and whose 1-skeleton $X^{1}$ has exactly one vertex on each boundary component of $\mathcal{S}(X)$. Among tubes, a biface is a portalgon $B$ of two triangles whose respective sides $s_{0},s_{1},s_{2}$ and $s_{0}^{\prime},s_{1}^{\prime},s_{2}^{\prime}$, in order (clockwise say), are such that $s_{0}$ is matched with $s_{0}^{\prime}$ and $s_{1}$ is matched with $s_{1}^{\prime}$. Its 1-skeleton $B^{1}$ has four edges: two loop edges forming the two boundary components of $\mathcal{S}(B)$, which we call boundary edges, and two edges whose relative interiors are included in the interior of $\mathcal{S}(B)$, which we call interior edges.
Figure 2: (From left to right) A good biface, a biface not good, a thin biface, a thick biface.
@@
Proposition 3.2 is similar to a result described by Löffler, Ophelders, Silveira, and Staals [portalgons, Theorem 45], building upon a ray shooting algorithm of Erickson and Nayyeri [erickson2012tracing].
@@ 4.1 Inserting vertices and edges
Appendix B details how to modify the portalgon $T$ to perform DeleteVertices. Applying InsertVertices to a triangular portalgon $T$ produces a portalgon $T^{\prime}$ whose polygons are usually not triangles. We now consider transforming $T^{\prime}$ into a triangular portalgon. To do that we repeatedly cut the polygons of $T^{\prime}$. We need a definition. In the plane consider a polygon $P$, two vertices $u\neq v$ of $P$, and the rectilinear segment $a$ between $u$ and $v$. If the relative interior of $a$ is included in the interior of $P$ then $a$ is called a vertex-to-vertex arc of $P$. It is easily seen that if $P$ has at least is not a triangle then $P$ has at least one vertex-to-vertex arc. Among the vertex-to-vertex arcs of $P$, the shortest ones are the shortcuts of $P$. We emphasize that we consider the shortest ones among all the vertex-to-vertex arcs, without fixing the endpoints, but the endpoints are chosen among the vertices of $P$. In a portalgon $T$ every polygon $P$ corresponds to a face $F$ of $T^{1}$, and every shortcut of $P$ corresponds to a path whose relative interior is included in $F$: we say of this path that it is a shortcut of $F$.

### v2

@@ 1 Introduction
In one of its simplest forms a triangulation is a finite collection of disjoint triangles in the Euclidean plane, together with a partial matching of the sides of the triangles such that any two matched sides have the same length (Figure 2.1). This simple representation appears under different names in the literature (intrinsic triangulation [29, 30], portalgon [22]). Cutting out the triangles from the plane and identifying the matched sides isometrically, respecting the orientations of the triangles, provides a (compact, orientable) polyhedral surface. This surface is closed if, in addition, it is connected and without boundary.
In this paper we consider the Delaunay triangulation of a closed polyhedral surface whose vertex set consists of the singularities (the points surrounded by an angle distinct from $2\pi$) of the surface (except for flat tori, see below). It is generically unique. Our main contribution is an algorithm to compute it from any triangulation of the surface, whose time complexity is polynomial in the number of triangles and in the logarithm of the aspect ratio of the input triangulation. Our second contribution is a lower bound showing that the dependency in the logarithm of the aspect ratio is unavoidable in our model of computation.
@@ 1.1 Related works
Polyhedral surfaces can also be obtained from meshes, flat triangles in $\mathbb{R}^{3}$ glued along their edges. Moreover, every mesh defines a triangulation of its surface. Yet triangulations are more general than meshes: most triangulations cannot be obtained from a mesh. Some recent algorithms advantageously operate on triangulations of polyhedral surfaces without reference to a mesh [28, 34, 19]. In this context the adjective “intrinsic” is sometimes placed before the name “triangulation” to make the distinction with the particular triangulations arising from a mesh. In the mathematical community, a prominent example is that of a translation surface [23, 35, 16, 14], which arises naturally in the study of billiards in rational polygons.
@@
Triangulations are so general that not all of them are suitable for computation, compared to meshes. Prominently, a fundamental problem on polyhedral surfaces is to compute the distance or a report a shortest path between two points. On a mesh, shortest paths can be computed in time polynomial in the number of triangles. For example, an algorithm of Mitchell, Mount, and Papadimitriou [25] propagates waves along the surface, starting from the source, in a discrete manner. See also Chen and Han [5]. On a generic triangulation however (not arising from a mesh), the number of times a shortest path visits a triangle is not bounded by any function of the number of triangles, as noted for example almost 20 years ago by Erickson [12]. Recently, Löffler, Ophelders, Staals, and Silveira [22] coined the term happiness of a triangulation, for the maximum number of times a shortest path visits a triangle. They adapted the single-source shortest paths algorithm of Mitchell, Mount, and Papadimitriou [25] from meshes to triangulations, whose time complexity now depends on the happiness of the triangulation (it is more efficient on triangulations of low happiness).
@@
This raises the problem of replacing any given triangulation by another triangulation of the same surface whose happiness is “low”. Among the many remeshing algorithms [15, 31, 27, 32], only few have been ported to the general context of intrinsic triangulations [29], and the only solution we are aware of that compares to our main result (Theorem 1.1 below), by Löffler, Ophelders, Staals, and Silveira [22, Section 5], is restricted to particular inputs whose surfaces are all homeomorphic to an annulus; we will use it as a black box. Importantly, the same authors also showed that Delaunay triangulations have bounded happiness [22, Section 4.2].
@@
Delaunay triangulations are classical objects of computational geometry [6, 13, 2], closely related to shortest paths. While mostly known in the plane, they generalize to closed polyhedral surfaces, see for example the depiction of Bobenko and Springborn [4]. To compute a Delaunay triangulation from an arbitrary intrinsic triangulation there are, to our knowledge, only two approaches, and neither compares to our main result (Theorem 1.1). One approach computes a Voronoi diagram with a suitably adapted multiple-source shortest path algorithm, and then derives from it a Delaunay tessellation, see for example Mount [26], Liu, Chen, and Tang [20], and Liu, Xu, Fan, and He [21]. Another approach starts from an initial triangulation and flips its edges until it reaches a Delaunay triangulation; it was proved to terminate by Indermitte, Liebling, Troyanov, and Clémençon [4, 17].
@@ 1.2 Our results
In order to state our results precisely, it now matters to make the distinction between a triangulation and the data structure representing it, and to allow for more general polygons than triangles. Following Löffler, Ophelders, Staals, and Silveira [22], we call portalgon the collection $T$ of polygons in the Euclidean plane and the partial matching of their sides. We denote by $\mathcal{S}(T)$ the associated polyhedral surface. We say that the portalgon $T$ is triangular if all polygons are triangles. The sides of the polygons, once identified, constitute a graph $T^{1}$ embedded on $\mathcal{S}(T)$ (the polygons themselves correspond to the faces of $T^{1}$): it is this graph $T^{1}$ that we call triangulation if $T$ is triangular, and we call $T^{1}$ a tessellation in general.
We consider, on a closed polyhedral surface $S$, the unique Delaunay tessellation $\mathcal{D}$ of $S$ whose vertices are exactly the singularities of $S$, with a single very special exception: if $S$ has no singularity, then $S$ is a flat torus and we let $\mathcal{D}$ be any of the Delaunay tessellations of $S$ that have exactly one vertex, for one can be mapped to the other via an orientation-preserving isometry of $S$ anyway. In any case, we say that $\mathcal{D}$ is the Delaunay tessellation of $S$, in a slight abuse. [^2] Given a set $V$ of points of the surface, finite, non-empty, and containing all the singularities, our results easily extend to Delaunay triangulations whose vertex set is $V$, but this is incidental to us. It is “generically” a triangulation, but not always. If not, then triangulating the faces of $\mathcal{D}$ along any vertex-to-vertex arcs provides a Delaunay triangulation. The aspect ratio of a triangular portalgon $T$ is the maximum side length of a triangle of $T$ divided by the smallest height of a triangle of $T$ (possibly another triangle). Our main contribution is:
@@
We analyze our algorithms within the real RAM model of computation described by Erickson, van der Hoog, and Miltzow [11]. It is an extension of the standard integer word RAM, with an additional memory array storing reals, and with additional instructions. On such a machine, we represent each polygon of a portalgon $T$ by the list of its vertices, and each vertex is by its two coordinates, stored in the memory array dedicated to reals. So displacing (translating or rotating) the polygons in the plane provides different representations of $T$. When modifying a portalgon $T$, we actually modify our representation of $T$, using elementary operations that are easily seen to be achievable by a real RAM.
@@
Proof of Theorem 1.1.
Proposition 1.4 computes in $O(n\log^{2}(n)\cdot\log^{2}(r))$ time a portalgon $T^{\prime}$ of $O(n\cdot\log(r))$ triangles, whose happiness is $O(n\log(n)\cdot\log^{2}(r))$. Proposition 1.5 then computes the portalgon of the Delaunay tessellation from $T^{\prime}$ in $O(n^{3}\log^{2}(n)\cdot\log^{4}(r))$ time. ∎
@@ 2 Preliminaries
We use without review standard notions of graph theory and low dimensional topology and geometry, referring to textbooks for details [7, 1, 33, 8]. We only mention that on a surface $S$, a path $p:[0,1]\to S$ is simple if its restriction to the interval $(0,1)$ is injective, in which case the image of $(0,1)$ by $p$ is the relative interior of $p$. We denote by $\ell(p)$ the length of a geodesic path $p$. Throughout the paper, logarithms are in base two.
The definition of Delaunay tessellation given by Bobenko and Springborn [4, Section 2] is not used in the core of the paper, but only in in Appendix J for proving Proposition 1.5. We collect details on this definition in Appendix I for completeness.
@@ 2.1 Portalgons, tessellations, and polyhedral surfaces
A portalgon $T$ is a disjoint collection of oriented polygons in the Euclidean plane, together with a partial matching of the sides of the polygons such that any two matched sides have the same length. It is triangular if all polygons are triangles. See Figure 2.1. Any subset of the polygons defines a sub-portalgon $T^{\prime}$ of $T$: two sides of polygons are matched in $T^{\prime}$ if and only if they are matched in $T$. In a portalgon $T$, identifying the matched sides, isometrically, and respecting the orientations of the polygons, provides the surface of $T$, denoted $\mathcal{S}(T)$; it is a 2-dimensional Riemannian manifold whose metric may have singularities. The sides of the polygons of $T$ correspond to a graph $T^{1}$ embedded on $\mathcal{S}(T)$, the 1-skeleton of $T$.
@@
Figure 2.1: (Left) A triangular portalgon $T$: two triangles in the Euclidean plane, with two sides matched in red. (Right) The surface $\mathcal{S}(T)$, and the 1-skeleton $T^{1}$.
@@ 2.2 Aspect ratio, systole, happiness, and segment-happiness
The aspect ratio of a triangular portalgon $T$ is the maximum side length of a triangle of $T$ divided by the smallest height of a triangle of $T$ (possibly another triangle). Note that the aspect ratio is always greater than or equal to $2/\sqrt{3}>1$, because the maximum side length of a triangle is always greater than or equal to $2/\sqrt{3}$ times its smallest height.
@@
The happiness of a portalgon $T$ is the maximum number of times a shortest path in $\mathcal{S}(T)$ visits the image of a polygon of $T$, maximized over all the shortest paths of $\mathcal{S}(T)$ and all the polygons of $T$ (see [22, Section 3]). We introduce a variation, more suitable to our needs. In a polyhedral surface $S$, a segment is a simple geodesic path $e$ whose relative interior is disjoint from any singularity of $S$. The segment-happiness of $e$ in $S$, denoted $h_{S}(e)$, is the maximum number of intersections between $e$ and a shortest path of $S$, maximized over all the shortest paths of $S$. The segment-happiness of a portalgon $T$ is then the maximum segment-happiness $h_{\mathcal{S}(T)}(e)$, maximized over the edges $e$ of its 1-skeleton $T^{1}$. A priori, the segment-happiness of a portalgon $T$ differs from the happiness of $T$. Indeed a path in $\mathcal{S}(T)$ may visit many times a face of $T^{1}$ without intersecting any edge of $T^{1}$ more than once, if the face has high degree. However, if $T$ is triangular, then the happiness and the segment-happiness of $T$ do not differ by more than a constant factor.
@@ 3 Tubes and bifaces
In this section we focus on particular triangular portalgons. See Figure 3.1. A tube is a triangular portalgon $X$ whose surface $\mathcal{S}(X)$ is homeomorphic to an annulus and has no singularity in its interior, and whose 1-skeleton $X^{1}$ has exactly one vertex on each boundary component of $\mathcal{S}(X)$. Among tubes, a biface is a portalgon $B$ of two triangles whose respective sides $s_{0},s_{1},s_{2}$ and $s_{0}^{\prime},s_{1}^{\prime},s_{2}^{\prime}$, in order (clockwise say), are such that $s_{0}$ is matched with $s_{0}^{\prime}$ and $s_{1}$ is matched with $s_{1}^{\prime}$. Its 1-skeleton $B^{1}$ has four edges: two loop edges forming the two boundary components of $\mathcal{S}(B)$, which we call boundary edges, and two edges whose relative interiors are included in the interior of $\mathcal{S}(B)$, which we call interior edges.
Figure 3.1: (From left to right) A good biface, a biface not good, a thin biface, a thick biface.
@@
Proposition 3.2 is similar to a result described by Löffler, Ophelders, Silveira, and Staals [22, Theorem 45], building upon a ray shooting algorithm of Erickson and Nayyeri [10].
@@ 4.1 Inserting vertices and edges
Appendix B details how to modify the portalgon $T$ to perform InsertVertices. Applying InsertVertices to a triangular portalgon $T$ produces a portalgon $T^{\prime}$ whose polygons are usually not triangles. We now consider transforming $T^{\prime}$ into a triangular portalgon. To do that we repeatedly cut the polygons of $T^{\prime}$. We need a definition. In the plane consider a polygon $P$, two vertices $u\neq v$ of $P$, and the rectilinear segment $a$ between $u$ and $v$. If the relative interior of $a$ is included in the interior of $P$ then $a$ is called a vertex-to-vertex arc of $P$. It is easily seen that if $P$ is not a triangle then $P$ has at least one vertex-to-vertex arc. Among the vertex-to-vertex arcs of $P$, the shortest ones are the shortcuts of $P$. We emphasize that we consider the shortest ones among all the vertex-to-vertex arcs, without fixing the endpoints, but the endpoints are chosen among the vertices of $P$. In a portalgon $T$ every polygon $P$ corresponds to a face $F$ of $T^{1}$, and every shortcut of $P$ corresponds to a path whose relative interior is included in $F$: we say of this path that it is a shortcut of $F$.